niyahissassy:

(12-3)x(10+20)x(14-6)= a.1,000 b.60 c,2160

supie:

Do the \(x\)'s represent multiplication or a variable?

niyahissassy:

do yu know

niyahissassy:

variable

supie:

@jhonyy9

ramen:

I'm almost 100 sure that those X's mean multiplication, they aren't variables.

supie:

Same...

ramen:

Because when I did the problem, I got one of the answer choices.

supie:

\((12−3)(10+20)(14−6)=\) \(9(10+20)(14−6)=\) \((9)(30)(14−6)=\) \(270(14−6)=\) \(270)(8)=?\)

niyahissassy:

...